OpenStudy (anonymous):

I've recently taken a math quiz, and there was one problem that wouldn't get out of my head: e^(2x+1)=6*e^(x+3) Solve for x. My answer was: x=ln(6)+2 I now know that's the wrong answer and a calculator states it is x=2. Could anyone perhaps tell me how to solve for x? In a step-by-step process; if another problem comes up in the future, I'd like to think back and remember all the steps.

OpenStudy (anonymous):

You are actually correct? http://www.wolframalpha.com/input/?i=e%5E%282x%2B1%29%3D6*e%5E%28x%2B3%29

OpenStudy (anonymous):

divide both sides by \[e^{x+3}\]

OpenStudy (anonymous):

get \[\frac{e^{2x+1}}{e^{x+3}}=6\]

OpenStudy (anonymous):

subtract the exponents , get \[e^{x-2}=6\] making \[x-2=\ln(6)\] or \[x=\ln(6)+2\]

OpenStudy (jdoe0001):

yeap.... got the same here as well

OpenStudy (anonymous):

if it was \[e^{2x+1}=e^{x+3}\] THEN it would be \(x=2\)

OpenStudy (jdoe0001):

I'm reading \(\Large \bf e^{2x+1}=6\cdot e^{x+3}\) if so, then ln(6)+2 is right

OpenStudy (anonymous):

Yay!, Mathway Calculator stated x=2, and my math quiz was 3 questions, I thought i was so screwed.

OpenStudy (anonymous):

Thank you, everyone.

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Here's my way, if you are interesting...: I substitute 6 to \(\large e^{\ln6}\) so we have: \(\large e^{x-2}=e^{\ln6}e^{x+3} \\~\\ \large \large e^{x-2}=e^{x+3+\ln6} \\~\\x-2=x+3+\ln6\) You can do the rest from here

OpenStudy (anonymous):

Oops i typed x instead of 2x in left side of equation... sorry

OpenStudy (anonymous):

actually 2x+1, not x-2 haha sorry about that... but you get the idea; replace 6 to \(e^{ln6}\)

OpenStudy (anonymous):

Thank you! I didn't realize there were so many ways to do it. On the quiz i started off by using the inverse ln: ln(e^(2x+1))=ln(6)*ln(e^(x+3)) =ln(e^2x+1))=ln(6+e^(x+3)) =(2x+1)=ln(6)+(x+3) and had continued from there.